Required information [The following information applies to the questions displayed below.] On October 29, Lobo Co. began operations by purchasing razors for resale. The razors have a 90-day warranty. When a razor is returned, the company discards it and mails a new one from Merchandise Inventory to the customer. The company’s cost per new razor is $20 and its retail selling price is $75. The company expects warranty costs to equal 8% of dollar sales. The following transactions occurred. Nov. 11 Sold 105 razors for $7,875 cash. 30 Recognized warranty expense related to November sales with an adjusting entry. Dec. 9 Replaced 15 razors that were returned under the warranty. 16 Sold 220 razors for $16,500 cash. 29 Replaced 30 razors that were returned under the warranty. 31 Recognized warranty expense related to December sales with an adjusting entry. Jan. 5 Sold 150 razors for $11,250 cash. 17 Replaced 50 razors that were returned under the warranty. 31 Recognized warranty expense related to January sales with an adjusting entry. 1. Prepare journal entries to record above transactions and adjustments.
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 7,875
Sales 7,875
11-Nov Cost of goods sold 2100 =105*20
Merchandise inventory 2,100
30-Nov Warranty expense 630 =7875*8%
Estimated warranty liability 630
9-Dec Estimated warranty liability 300 =15*20
Merchandise inventory 300
16-Dec Cash 16,500
Sales 16,500
16-Dec Cost of goods sold 4,400 =220*20
Merchandise inventory 4,400
29-Dec Estimated warranty liability 600 =30*20
Merchandise inventory 600
31-Dec Warranty expense 1,320 =16500*8%
Estimated warranty liability 1,320
1.2
Date General Journal Debit Credit
5-Jan Cash 11,250
Sales 11,250
5-Jan Cost of goods sold 3,000 =150*20
Merchandise inventory 3,000
17-Jan Estimated warranty liability 1000 =50*20
Merchandise inventory 1,000
31-Jan Warranty expense 900 =11250*8%
Estimated warranty liability 900